OpenStudy (anonymous):

solution for 2x²=x ?

OpenStudy (anonymous):

\[2x²=x \rightarrow \frac{x^2}{x} = \frac{1}{2} \rightarrow x = \frac{1}{2}\] @professor.obakeng

OpenStudy (phi):

you can write it in standard form \[ 2x^2-x=0 \] factor it to get \[ x(2x-1)= 0 \] you get two solutions: either x=0 or 2x-1=0 which means 2x=1 or x=½